Suppose sinA = 12/13 with 90º≤A≤180º. Suppose also that sinB = -7/25 with -90º≤B≤0º. Find cos(A - B).
Answer by lwsshak3(11628)

Suppose sinA = 12/13 with 90º≤A≤180º. Suppose also that sinB = -7/25 with -90º≤B≤0º. Find cos(A - B).
**
cos addition formula: Cos(A-B)=cosAcosB+sinAsinB
given sin A=12/13=opposite side/hypotenuse
adjacent side=√(13^2-12^2)=√(169-144)=√25=5
cos A=adjacent side/hypotenuse=-5/13 (in quadrant II where cos<0)
..
given sin B=-7/25=opposite side/hypotenuse
adjacent side=√(25^2-7^2)=√(625-49)=√576=24
cos B=adjacent side/hypotenuse=24/25 (in quadrant IV where cos>0)
..
Cos(A-B)=-5/13*24/25+12/13*+7/25=-120/325+84/325=-36/325≈-.1108